OpenStudy (anonymous):

A right triangle has legs of lengths 8 and 15. How do you find the hypotenuse?

OpenStudy (anonymous):

pythagoras thoeram. . .

OpenStudy (anonymous):

a^2+b^2=c^2?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

use pythagoras rule i.e a^+b^2=c^2 where a=perpendicular b=base c=hypotenuse

OpenStudy (anonymous):

23^2

OpenStudy (anonymous):

17 is the answer

OpenStudy (anonymous):

@beezz12 yup it is the right answer